OpenStudy (anonymous):

if v =+/-(x-1) however when x=0m, v= 1m/s >0 why do we take v = -(x-1)

OpenStudy (ash2326):

@virtus do you have a graph along with this?

OpenStudy (anonymous):

nope i do not have a graph sorry

OpenStudy (ash2326):

@virtus Is this the full question?

OpenStudy (anonymous):

a particle moves along a straight line s that its acceleration is given by a =x-1 where x is its displacement from the origin. Initially, the particles ias at the origin and has velocity of 1m/s

OpenStudy (anonymous):

find x as a function of t and describe the motion of the particles

OpenStudy (ash2326):

OK, so we have \[a=x-1\] a=acceleration we are given that at t=0, velocity=1m/s , x=0 Do you agree with this?

OpenStudy (anonymous):

@ash2326 I dont think x is always 0 when t=0

OpenStudy (ash2326):

Here it's given that "Initially, the particles ias at the origin and has velocity of 1m/s"

OpenStudy (anonymous):

oh ya....

OpenStudy (ash2326):

@virtus ??

OpenStudy (anonymous):

ok....

OpenStudy (ash2326):

We know that acceleration a \[a=\frac{d^2x}{dt^2}\] so we get \[\frac{d^2x}{dt^2}=x-1\]

OpenStudy (ash2326):

Do you understand this?

OpenStudy (anonymous):

yeah

OpenStudy (ash2326):

It's better we use differential equations to solve this Do you know differential equations?

OpenStudy (ash2326):

@virtus ??

OpenStudy (anonymous):

no

OpenStudy (ash2326):

we have \[a=x-1\] we know \[a=\frac {dv}{dt}\] so \[\frac{dv}{dt}=x-1\] where v = velocity But here the function is in terms of x, so we'll have to change variables we know that \(v=\frac {dx}{dt}\) \[\frac{dv}{dt}=x-1\] multiplying and dividing left side by dx we get \[\frac{dv\times dx}{dt\times dx}=x-1\] or \[\frac{dv}{dx}\times\frac {dx}{dt}=x-1\] we know that \(v=\frac {dx}{dt}\) so we get \[\frac{dv}{dx}\times v=x-1\] do you get these steps @virtus ??

OpenStudy (anonymous):

thank you i understand

OpenStudy (ash2326):

it's not over yet, still we have some work. Are you here?